OpenStudy (anonymous):

Find the sum. one-twelfth plus three-fourths one-third seven-twelfths two-thirds five-sixths

OpenStudy (anonymous):

5/6

OpenStudy (diyadiya):

\[\frac{1}{12}+ \frac{3}{4}\]Taking LCM \[\frac{1}{12}+ \frac{3 \times 3}{4 \times 3}= \frac{1}{12}+ \frac{9}{12}= \frac{10}{12}= \frac{5}{6}\]